# 3M-TI: High-Quality Mobile Thermal Imaging via Calibration-free Multi-Camera Cross-Modal Diffusion

### v1

@@ Abstract
The miniaturization of thermal sensors for mobile platforms inherently limits their spatial resolution and textural fidelity, leading to blurry and less informative images. Existing thermal super-resolution (SR) methods can be grouped into single-image and RGB-guided approaches: the former struggles to recover fine structures from limited information, while the latter relies on accurate and laborious cross-camera calibration, which hinders practical deployment and robustness. Here, we propose 3M-TI, a calibration-free Multi-camera cross-Modality diffusion framework for Mobile Thermal Imaging. At its core, 3M-TI integrates a cross-modal self-attention module (CSM) into the diffusion UNet, replacing the original self-attention layers to adaptively align thermal and RGB features throughout the denoising process, without requiring explicit camera calibration. This design enables the diffusion network to leverage its generative prior to enhance spatial resolution, structural fidelity, and texture detail in the super-resolved thermal images. Extensive evaluations on real-world mobile thermal cameras and public benchmarks validate our superior performance, achieving state-of-the-art results in both visual quality and quantitative metrics. More importantly, the thermal images enhanced by 3M-TI lead to substantial gains in critical downstream tasks like object detection and segmentation, underscoring its practical value for robust mobile thermal perception systems. More materials: https://github.com/work-submit/3MTI.
@@ 4.1 Experimental Setup
Smartphone Dataset. For real-world evaluation, we collected 40 image pairs across 30 diverse scenes using our smartphone-based multi-camera system, including 37 nighttime and the remaining daytime captures. Similarly, all thermal images are resized to $64\times 64$. The corresponding RGB images are center-cropped to a square aspect ratio and resized to $512\times 512$ as high-resolution references.
@@
Comparative Methods. We compare our model against 5 representative baselines: CoReFusion [17], SwinFuSR [1], SwinPaste [63], SeeSR [52], and OSEDiff [51]. CoReFusion is an RGB-guided super-resolution model built upon the UNet architecture. SwinFuSR employs a Swin Transformer backbone with RGB guidance, while SwinPaste is an enhanced variant. SeeSR and OSEDiff are diffusion-based image super-resolution methods without reference guidance. For fair comparison, all baseline models are retrained on our dataset using their publicly released codes.
@@ 4.2 Image Quality Assesment
Table 1 summarizes the quantitative results. Our 3M-TI achieves the best performance on perceptual metrics, including LPIPS, MANIQA, and MUSIQ. For the reference-based LPIPS metric, 3M-TI, OSEDiff, and SeeSR outperform non-diffusion methods (CoReFusion, SwinFuSR, and SwinPaste), highlighting the power of diffusion priors for perceptual enhancement. In terms of fidelity metrics (PSNR, SSIM), 3M-TI ranks a close second, demonstrating better structural preservation than OSEDiff, SeeSR, SwinFuSR, and SwinPaste. The proposed misalignment augmentation significantly improves 3M-TI, particularly on perceptual metrics, while other methods see only marginal gains. This indicates that 3M-TI effectively learns cross-modal correspondences and generalizes well to diverse uncalibrated conditions.
@@ 4.3 Downstream Applications
Ours w/ All 30.09 0.8610 0.1787 36.66
@@ 4.4 Ablation Study
We conduct an ablation study to investigate the contributions of each component. Table 4 and Fig. 7 report quantitative drops and representative visual examples when individual modules are removed. Removing the RGB reference produces markedly blurrier reconstructions (e.g., bicycle spokes and shrub textures), indicating the importance of cross-modal cues. Removing the misalignment augmentation reduces robustness to geometric and temporal offsets, leading to visibly degraded high-frequency details. Removing the skip connection degrades structural fidelity: circular wheels become distorted and geometric consistency is lost. With all components enabled, 3M-TI attains the best balance of fidelity and perceptual quality, producing sharp, semantically consistent thermal reconstructions.

### v2

@@ Abstract
The miniaturization of thermal sensors for mobile platforms inherently limits their spatial resolution and textural fidelity, leading to blurry and less informative images. Existing thermal super-resolution (SR) methods can be grouped into single-image and RGB-guided approaches: the former struggles to recover fine structures from limited information, while the latter relies on accurate and laborious cross-camera calibration, which hinders practical deployment and robustness. Here, we propose 3M-TI, a calibration-free Multi-camera cross-Modality diffusion framework for Mobile Thermal Imaging. At its core, 3M-TI integrates a cross-modal self-attention module (CSM) into the diffusion UNet, replacing the original self-attention layers to adaptively align thermal and RGB features throughout the denoising process, without requiring explicit camera calibration. This design enables the diffusion network to leverage its generative prior to enhance spatial resolution, structural fidelity, and texture detail in the super-resolved thermal images. Extensive evaluations on real-world mobile thermal cameras and public benchmarks validate our superior performance, achieving state-of-the-art results in both visual quality and quantitative metrics. More importantly, the thermal images enhanced by 3M-TI lead to substantial gains in critical downstream tasks like object detection and segmentation, underscoring its practical value for robust mobile thermal perception systems. Project page: https://github.com/work-submit/3MTI
00footnotetext: * These authors have contributed equally to this work.00footnotetext: Corresponding author. Mail: yuanxiaoyun@sjtu.edu.cn.
@@ 4.1 Experimental Setup
Smartphone Dataset. For real-world evaluation, we collected 300 image pairs across 100 diverse scenes using our smartphone-based multi-camera system, including 200 nighttime and the remaining daytime captures. Similarly, all thermal images are resized to $64\times 64$. The corresponding RGB images are center-cropped to a square aspect ratio and resized to $512\times 512$ as high-resolution references.
@@
Comparative Methods. We compare our model against 7 representative baselines: CoReFusion [17], CoRPLE [23], SwinFuSR [1], SwinPaste [64], SeeSR [53], OSEDiff [52], and DifIISR [24]. CoReFusion is an RGB-guided super-resolution model built upon the UNet architecture. SwinFuSR employs a Swin Transformer backbone with RGB guidance, while SwinPaste is an enhanced variant. SeeSR, OSEDiff, and DifIISR are diffusion-based image super-resolution methods without reference guidance. Among them, DifIISR is specifically designed for infrared images. CoRPLE is a hybrid framework combining contourlet residual and prompt learning for infrared image enhancement. For fair comparison, all baseline models except DifIISR are retrained on our dataset using their publicly released codes, while DifIISR is evaluated directly since its training code is unavailable.
@@ 4.2 Image Quality Assesment
Table 1 summarizes the quantitative results. Our 3M-TI achieves the best performance on perceptual metrics, including LPIPS, MANIQA, and MUSIQ. For the reference-based LPIPS metric, 3M-TI, OSEDiff, and SeeSR outperform non-diffusion methods (CoReFusion, CoRPLE, SwinFuSR, and SwinPaste), highlighting the power of diffusion priors for perceptual enhancement. In terms of fidelity metrics (PSNR, SSIM), 3M-TI demonstrate better structural preservation than OSEDiff, SeeSR, DifIISR, SwinFuSR, and SwinPaste. The proposed misalignment augmentation significantly improves 3M-TI, particularly on perceptual metrics, while other methods see only marginal gains. This indicates that 3M-TI effectively learns cross-modal correspondences and generalizes well to diverse uncalibrated conditions.
@@ 4.3 Downstream Applications
w/ Self-Attn 30.03 0.8591 0.1954 34.47
w/ Feature Concat 29.60 0.8481 0.2164 32.12
w/ Cross-Attn 29.86 0.8555 0.1953 34.32
w/ All (w/ CSM) 30.09 0.8610 0.1787 36.66
@@ 4.4 Ablation Study
We conduct an ablation study to investigate the contributions of each component. Table 4 and Fig. S5 in the supplementary material report quantitative drops and representative visual examples when individual modules are removed. Removing the RGB reference produces markedly blurrier reconstructions (e.g., bicycle spokes and shrub textures), indicating the importance of cross-modal cues. Removing the misalignment augmentation reduces robustness to geometric and temporal offsets, leading to visibly degraded high-frequency details. Removing the skip connection degrades structural fidelity: circular wheels become distorted and geometric consistency is lost. Removing the semantic guidance of RAM hardly impairs reference metrics yet brings a drop in perceptual quality. With all components enabled (w/ All), 3M-TI attains the best balance of fidelity and perceptual quality, producing sharp, semantically consistent thermal reconstructions.
We further compared CSM (w/ CSM) against: 1) Original Self-Attn (intra-modal attention only); 2) Feature Concat + FC (static projection w/o attention); 3) Standard Cross-Attn (inter-modal attention only). The lower half of Table 4 show that CSM outperforms all variants. Original Self-Attn is incapable of fusing RGB cues. Feature Concat relies on static weights, lacking content-adaptive capabilities. Standard Cross-Attn neglects the internal spatial context within the thermal modality itself. In contrast, CSM concatenates tokens to perform joint Self-Attn. This enables the model to capture both inter-modal (RGB-thermal) guidance and intra-modal (thermal-thermal) structural dependencies, which is vital for integrating complementary RGB details while preserving thermal structures.
@@ 5 Conclusion
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 62271283, and in part by the Shanghai Jiao Tong University Medical-Engineering Cross Research Fund under Grant No. YG2026QNB50.